# Multimodal Causal Reasoning Benchmark: Challenging Multimodal Large Language Models to Discern Causal Links Across Modalities

### v3

@@ A.4 Producing Contextual Description Pairs
The absence of crucial visual cues could introduce randomness in image creation, which may lead to inconsistencies and potentially undermine the perceived causality between the siamese images. Recent research on content consistency Chen et al. (2025); Long et al. (2024); long2024video has become popular in long video generation by maintaining coherent content across frames. For image content consistency, Figure 18 highlights the drawbacks of missing context and the advantages of incorporating context. As shown in Figure 18 (a), although the two columns of images accurately represent the core caption, mismatched clothing disrupts the sense of causality, making it difficult to form coherent pairs. In contrast, the example in Figure 18 (b) demonstrates that incorporating contextual information and transforming core captions into contextual descriptions effectively resolves this issue and reduces randomness in image synthesis. To achieve this, we leverage the linguistic capabilities of LLMs to enhance core caption pairs by enriching contextual details such as appearance, clothing, environment, and atmosphere. Additionally, we introduce subtle changes, such as variations in facial expressions, within the contextual description pairs to reflect the passage of time. These detailed variations emphasize the impact of causality over time, making the connection between siamese images more natural and coherent.
@@
we also compare identity-preserving techniques with our prompt-guidance method (Figure 19). Traditional identity-preserving image synthesis methods (e.g., LCM Gal et al. (2024) and IP-Adapter Ye et al. (2023)) focus on image personalization by retaining identity details through a region encoder during the generation process Wang et al. (2024b). However, this approach leads to two major issues. First, most existing identity-preserving techniques rely heavily on guided images, which limits their capacity for semantically-driven image generation and requires finding a suitable ID image for each causal scenario. Second, as the name suggests, identity-preserving methods focus primarily on maintaining facial identity (appearance) but struggle to incorporate cause-and-effect relationships across images. In contrast, our causal-and-effect image synthesis approach leverages the linguistic capabilities of large language models (LLMs) to integrate a richer spectrum of contextual information. It not only preserves human facial identity (appearance), but also accounts for additional details (e.g., clothing, environment, and overall atmosphere). This ensures that images remain coherent even when modifications are introduced through causal reasoning.
@@ C.2 Form of Visual Input
Our case analysis demonstrates that, compared to Form-3, Forms-1 and Form-2 impose limitations on MLLMs’ ability to recognize and leverage critical visual cues necessary for multimodal causal reasoning. As shown in Figure 24, Form-3 provides GPT-4o with direct visual information, enabling it to successfully identify essential details, such as the continuity in a person’s appearance across cause-and-effect images. This was evident in GPT-4o’s output, where it correctly determined that the woman in the cause image, overwhelmed by paperwork, was the same individual in the effect image, now engaged in a serious discussion about work. This recognition of visual consistency is crucial for establishing causal relationships. However, when using Form-1, GPT-4o was unable to incorporate this specific visual cue and instead selected a different effect image (a generic team meeting) based on a more abstract textual interpretation rather than a direct visual correlation.
The key issue with Forms-1 and Form-2 is that they rely on structured textual descriptions that predefine categories of reasoning, which may inadvertently filter out implicit but important visual details. These formats encourage MLLMs to focus on generalized textual patterns rather than independently deriving causal relationships from visual features like facial expressions, body language, and scene continuity. By contrast, Form-3 allows GPT-4o to analyze raw visual inputs more freely, enhancing its ability to establish causal links based on direct visual observation rather than abstracted textual hints. This distinction highlights the potential shortcomings of rigid textual input structures in multimodal causal reasoning tasks. While textual guidance can be helpful, it may also constrain the model’s reasoning process, making it less sensitive to nuanced visual cues. Ensuring that MLLMs receive input formats that preserve rich visual information is therefore essential for improving their ability to perform causal inference in multimodal settings.

### v4

@@ A.4 Producing Contextual Description Pairs
The absence of crucial visual cues could introduce randomness in image creation, which may lead to inconsistencies and potentially undermine the perceived causality between the siamese images. Recent research on content consistency Chen et al. (2025); Long et al. (2024) has become popular in long video generation by maintaining coherent content across frames. For image content consistency, Figure 18 highlights the drawbacks of missing context and the advantages of incorporating context. As shown in Figure 18 (a), although the two columns of images accurately represent the core caption, mismatched clothing disrupts the sense of causality, making it difficult to form coherent pairs. In contrast, the example in Figure 18 (b) demonstrates that incorporating contextual information and transforming core captions into contextual descriptions effectively resolves this issue and reduces randomness in image synthesis. To achieve this, we leverage the linguistic capabilities of LLMs to enhance core caption pairs by enriching contextual details such as appearance, clothing, environment, and atmosphere. Additionally, we introduce subtle changes, such as variations in facial expressions, within the contextual description pairs to reflect the passage of time. These detailed variations emphasize the impact of causality over time, making the connection between siamese images more natural and coherent.
@@
We also compare identity-preserving techniques with our prompt-guidance method (Figure 19). Traditional identity-preserving image synthesis methods (e.g., LCM Gal et al. (2024) and IP-Adapter Ye et al. (2023)) focus on image personalization by retaining identity details through a region encoder during the generation process Wang et al. (2024b). However, this approach leads to two major issues. First, most existing identity-preserving techniques rely heavily on guided images, which limits their capacity for semantically-driven image generation and requires finding a suitable ID image for each causal scenario. Second, as the name suggests, identity-preserving methods focus primarily on maintaining facial identity (appearance) but struggle to incorporate cause-and-effect relationships across images. In contrast, our causal-and-effect image synthesis approach leverages the linguistic capabilities of large language models (LLMs) to integrate a richer spectrum of contextual information. It not only preserves human facial identity (appearance), but also accounts for additional details (e.g., clothing, environment, and overall atmosphere). This ensures that images remain coherent even when modifications are introduced through causal reasoning. The producing contextual description pairs prompt is organised as: "Task Overview: You need to convert causal-relevant captions into detailed image descriptions for generating images. Ensure the following: (A) Consistency in Characters: Use one sentence to describe the person’s appearance in each image description. Make sure two sentences in the two descriptions share the same information without using words like same or similar. (B) Face expression in Characters: If the descriptions contain the person’s appearance, please directly add a sentence following to describe the person’s facial expression, matching the scene. (C) Activities or Behaviors: Cause descriptions should exclusively detail the causal activities or behaviors, while effect descriptions should exclusively detail the resultant activities or behaviors. (D) Consistency in Scenes: If the scene remains unchanged between cause and effect, ensure the background description is identical in both descriptions. (E) Clear Causal Link: (1) Enhance Cause: Provide concrete details about what led to the effect. (2) Improve Effect: Ensure the effect is both visually and logically linked to the cause, using relatable or observable descriptions. "
@@ C.2 Form of Visual Input
Our case analysis demonstrates that, compared to Form-3, Form-1 and Form-2 impose limitations on MLLMs’ ability to recognize and leverage critical visual cues necessary for multimodal causal reasoning. As shown in Figure 24, Form-3 provides GPT-4o with direct visual information, enabling it to successfully identify essential details, such as the continuity in a person’s appearance across cause-and-effect images. This was evident in GPT-4o’s output, where it correctly determined that the woman in the cause image, overwhelmed by paperwork, was the same individual in the effect image, now engaged in a serious discussion about work. This recognition of visual consistency is crucial for establishing causal relationships. However, when using Form-1, GPT-4o was unable to incorporate this specific visual cue and instead selected a different effect image (a generic team meeting) based on a more abstract textual interpretation rather than a direct visual correlation.
The key issue with Form-1 and Form-2 is that they rely on structured textual descriptions that predefine categories of reasoning, which may inadvertently filter out implicit but important visual details. These formats encourage MLLMs to focus on generalized textual patterns rather than independently deriving causal relationships from visual features like facial expressions, body language, and scene continuity. By contrast, Form-3 allows GPT-4o to analyze raw visual inputs more freely, enhancing its ability to establish causal links based on direct visual observation rather than abstracted textual hints. This distinction highlights the potential shortcomings of rigid textual input structures in multimodal causal reasoning tasks. While textual guidance can be helpful, it may also constrain the model’s reasoning process, making it less sensitive to nuanced visual cues. Ensuring that MLLMs receive input formats that preserve rich visual information is therefore essential for improving their ability to perform causal inference in multimodal settings.